A school principle claims that the students in his school are above average intelligence. A random sample of 30 studets IQ scores have mean of 112.5.Is there sufficient evidence to support the principals claim?The mean population IQ is 100 with standard deviation of 1.5


1
Expert's answer
2022-01-13T09:22:08-0500

Let's find z-value:


"z = \dfrac{x-x_{mean}}{\sigma/\sqrt{n}}"

where "\sigma = 1.5, n=30, x=112.5, x_{mean}=100". Thus, obtain:


"z = \dfrac{112.5-100}{1.5/\sqrt{30}} \approx 45.6"

This value is much greater than the critical values of "z", thus, the evidence is sufficient to support the principals claim.


Answer. The evidence is sufficient to support the principals claim.